OpenStudy (anonymous):

If the equation of motion of a particle is given by s = B cos(ωt + δ), the particle is said to undergo simple harmonic motion. (a) Find the velocity of the particle at time t. v(t) = -Bω*sin(ωt+δ) (b) When is the velocity 0? ??? ***I only need help for part (b)! thanks :)

OpenStudy (anonymous):

velocity is zero when s= is said to be maximum so s= max when v=0=-Bwsin(wt+d) 0=sin(wt+d) arc sin 0=0=(wt+d) therefore -d=wt -d/w=t = time

OpenStudy (anonymous):

these are the only answer choices: http://img838.imageshack.us/img838/8430/capturebr.png

OpenStudy (anonymous):

so the velocity is zero at t=-d/w ans

OpenStudy (anonymous):

but that isnt one of the answer choices! :(

OpenStudy (anonymous):

can you give us the answer choices?

OpenStudy (anonymous):

it is linked above, in the imageshack link

OpenStudy (anonymous):

hmm i cant open the web site you given us......

OpenStudy (anonymous):

type them out here......

OpenStudy (anonymous):

It's the second one

OpenStudy (anonymous):

what did you see in the site heromiles? i cant open the site ..lol

OpenStudy (anonymous):

t = (npi - delta)/w

OpenStudy (anonymous):

to generalized this use n instead of 2 sin[npi-(wt+d)]=0 arc sin0=0= npi-(wt+d) wt+d=npi wt=npi-d t=(npi-d)/w ans

OpenStudy (anonymous):

so the velocity is zero at t=-d/w ans or in general: t=(npi-d)/w ans

OpenStudy (anonymous):

hope that understand you much better..lol

OpenStudy (anonymous):

Good job mark, but the OP is gone. He already submitted his test. So, I gave you a medal

OpenStudy (anonymous):

ah really? hahaha that was an actual test hehehe..lol thnx heromiles lol...have a lot of fun solving lol...

OpenStudy (anonymous):

I explained to him how easy it was already. All he had to do was set v'(t) to zero and solve for t.

OpenStudy (anonymous):

owkie thnx guys and enjoy..lol...